### v1

# One Look is Enough: A Novel Seamless Patchwise Refinement for Zero-Shot Monocular Depth Estimation Models on High-Resolution Images

Abstract
Zero-shot depth estimation (DE) models exhibit strong generalization performance as they are trained on large-scale datasets. However, existing models struggle with high-resolution images due to the discrepancy in image resolutions of training (with smaller resolutions) and inference (for high resolutions). Processing them at full resolution leads to decreased estimation accuracy on depth with tremendous memory consumption, while downsampling to the training resolution results in blurred edges in the estimated depth images. Prevailing high-resolution depth estimation methods adopt a patch-based approach, which introduces depth discontinuity issues when reassembling the estimated depth patches and results in test-time inefficiency. Additionally, to obtain fine-grained depth details, these methods rely on synthetic datasets due to the real-world sparse ground truth depth, leading to poor generalizability. To tackle these limitations, we propose Patch Refine Once (PRO), an efficient and generalizable tile-based framework. Our PRO consists of two key components: (i) Grouped Patch Consistency Training that enhances test-time efficiency while mitigating the depth discontinuity problem by jointly processing four overlapping patches and enforcing a consistency loss on their overlapping regions within a single backpropagation step, and (ii) Bias Free Masking that prevents the DE models from overfitting to dataset-specific biases, enabling better generalization to real-world datasets even after training on synthetic data. Zero-shot evaluation on Booster, ETH3D, Middlebury 2014, and NuScenes demonstrates into which our PRO can be well harmonized, making their DE capabilities still effective for the grid input of high-resolution images with little depth discontinuities at the grid boundaries. Our PRO runs fast at inference time.
{strip}[Uncaptioned image]
Figure 1: Qualitative comparison of patch-based depth estimation models (PatchFusion [23], PatchRefiner [24] and PRO (Ours)) for high-resolution images. PatchRefiner [24] and PRO (ours) also visualize the residuals to effectively highlight depth discontinuity issues in the 4th and 6th columns. All models performed depth estimations on $4\times 4$ patches of the 1st and 3rd input images. Zoom-in can help distinguish the grid lines (depth discontinuities) of the $4\times 4$ patches. Our proposed PRO achieves smooth boundary transitions without depth discontinuity artifacts along the grid boundaries.
1 Introduction
Monocular Depth Estimation (MDE) [10, 33, 36, 27, 43, 13, 14] has been widely investigated as the demand for 3D information continues to grow in autonomous driving, robotics, and virtual reality applications. After Midas [30] proposed an MDE network trainable with mixed training dataset, numerous zero-shot MDE networks [31, 9, 2, 46, 47, 19, 11] have been studied to predict depth for unseen real-world images. However, due to the architectural limitation or limited image resolutions of training dataset, most zero-shot MDE networks are trained specifically with the images of low resolutions (e.g. 384$\times$384, 518$\times$518). Consequently, when high resolution images are inputted to these models, the resulting estimated depth images tend to contain disrupted overall structures with low-frequency artifacts, although yielding improved edge details [26, 7].
Patch-based methods [26, 24, 23] have shown promising results by splitting high-resolution images into patches for DE, mitigating memory consumption issues while achieving remarkable performance. However, they suffer from depth discontinuity problems (e.g. boundary artifacts along grid boundaries) which occur when independent DE patches are reassembled to construct a complete (whole) depth map because depth continuity is maintained within each patch but not between patches. Previous methods [24, 23] alleviate this depth discontinuity problem (i) by incorporating a consistency loss during training and (ii) by ensemble averaging at test time which slows down inference speeds, making it impractical for real-world applications.
Figure 2: Performance and efficiency comparison on Middlebury 2014 [34]. The area of each circle represents the inference time. Circles of the same color represent the same model with different patch numbers ($\mathrm{P}=16$ for small-sized yellow and grey circles, $\mathrm{P}=177$ for large-sized yellow and grey circles). Our model (PRO) achieves the best performance in terms of edge errors ($\mathrm{D}^{3}\mathrm{R}$) and depth errors (AbsRel) while maintaining the fastest inference time. Inference time for all models is measured on the RTX 4090.
A common strategy for training high-resolution depth estimation is to train models on real-world high-resolution datasets [17, 20, 39, 42, 34] or leverage high-resolution synthetic datasets [38, 41, 40, 18]. Real-world datasets have the advantage of a smaller domain gap compared to synthetic datasets, but their ground truth (GT) depths are often sparse, particularly around edges, where supervision is crucial. Additionally, acquiring real-world dense depth data is challenging, limiting the DE learning for high-resolution images. On the other hand, synthetic datasets provide dense GT depth with fine details, but models trained on synthetic datasets often struggle with zero-shot inference on real-world data due to the domain gap [21]. Moreover, we observe that transparent objects do not have accurate depth annotations in the UnrealStereo4K dataset [38] that is the only synthetic dataset with 4K depth labeling, as they are annotated with the depth of the background behind the transparent objects.
To address the aforementioned problems, we propose Patch Refine Once (PRO), a novel refinement model that enables DE models to generate accurate DE results via only a single refinement per patch during inference (i.e.,“One Look”). Our PRO consists of two strategies: Grouped Patch Consistency Training and Bias Free Masking. Fig. 1 shows qualitative comparison of patch-based depth estimation models for high-resolution images. As shown, our PRO yields depth refinement results with little depth discontinuities along the grid boundaries while the very recent two state-of-the-art (SOTA) models exhibit depth discontinuity artifacts. Fig. 2 illustrates performance and efficiency comparison for SOTA depth refinement models and our PRO. As observed, our PRO provides the best performance in terms of $\mathrm{D}^{3}\mathrm{R}$, AbsRel and inference speed. Our contribution is summarized as:
  • We propose Grouped Patch Consistency Training (GPCT) strategy that can be well harmonized with existing DE models to mitigate the depth discontinuity problem for patch-wise DE approaches on high-resolution images. The GPCT does not require test-time ensemble, allowing a 12$\times$ faster inference speed compared to the SOTA patch-based method;
  • We introduce Bias Free Masking (BFM) to selectively apply supervision signals by masking out unreliable regions in images. The unreliable regions are the regions that have inaccurately annotated depth in GT, which is often observed in the UnrealStereo4K dataset. This is problematic without our BFM that efficiently prevents depth refinement models from overfitting to domain specific bias;
  • Our PRO achieves SOTA performance in zero-shot evaluation on Booster [29], ETH3D [35], Middlebury 2014 [34], and NuScenes [4] across all metrics, and runs efficiently at inference time (12$\times$ faster than PatchRefiner with 177 patches), compared to the very recent depth refinement methods for high-resolution images.
2 Related work
2.1 Zero-shot Monocular Depth Estimation
Early works on MDE models [1, 25, 10, 33, 49] focused on training on each dataset [12, 37], resulting in high performance on the trained dataset with degenerate performance on unseen datasets because of domain gap between datasets. Zero-shot MDE models that perform well on unseen (in-the-wild) images have been widely studied to improve generalization. MegaDepth [22] and DiverseDepth [48] construct large-sized datasets by collecting images from the Internet, improving adaptability on images from diverse scenes and conditions. They encourage efforts to scale up datasets to enhance zero-shot performance.
MiDaS [30] proposes a scale and shift invariant loss, which enables MDE networks to be trained on mixed diverse dataset, making the model robust to unseen datasets. DPT [31], Omnidata [9], and MiDaS v3.1 [2] improve DE performance by applying transformer architectures to replace CNN architectures. DepthAnything [46] introduces semi-supervised learning that utilizes 62M unlabeled images as pseudo labels. A dramatic increase in dataset size has demonstrated robust generalization performance. DepthAnythingv2 [47] points out that real-world depth datasets contain label noise and lack fine details. It is trained with pseudo GT labels obtained from a teacher model trained on a synthetic dataset, achieving high-quality depth maps.
Instead of scaling up datasets, some studies deploy prior knowledge from Diffusion Models for DE. Marigold [19] fine-tunes a pretrained Stable Diffusion [32] model on a relatively small synthetic dataset, showing competitive results. Geowizard [11] jointly estimates depth and surface normals by utilizing self-attention across different domains to enhance geometric consistency. However, zero-shot MDE models are trained on datasets of low resolutions (e.g., 384$\times$384, 518$\times$518), which results in degenerate performance on high resolution images. When high-resolution images are downsampled to the training resolution for inference, fine details are lost. On the other hand, performing inference with a larger resolution input preserves details but degrades depth accuracy.
2.2 High-Resolution Depth Estimation
High-resolution DE models aim to predict accurate depth while preserving fine details. SMD-Net [38] utilizes an implicit function [5] to predict mixture density for precise DE at object boundaries. Dai et al. [7] introduce a Poisson fusion-based depth map optimization method by applying guided filtering in a self-supervised learning framework. SDDR [21] proposes a self-distilled depth refinement method that generates pseudo edge labels, addressing local consistency issues and edge deformation noise with Noisy Poisson Fusion. Patch-based high-resolution DE methods select patches and merge patch-wise results to improve depth details. BoostingDepth [26] introduces a patch selection process based on edge density, and employs iterative refinement through multi-resolution depth merging. PatchFusion [23] eliminates the need for patch selection by splitting patches according to a predefined method and utilizes shifted window self-attention to fuse global and local information. PatchRefiner [24] initializes depth estimation with a low-resolution depth map and predicts residuals, and proposes a detail and disentangling (DSD) loss for training on both real-world and synthetic datasets. However, patch-based methods inherently suffer from depth discontinuities at patch boundaries, as they emphasize local information over global consistency, or maintain depth continuity only within individual patches. To address this problem, prior works [23, 24] employ a significantly large number of patches for test-time ensemble, which considerably slows down inference speed. By contrast, our PRO (Patch Refine Once) model achieves fine-grained details efficiently by performing only a single refinement per patch during inference, despite utilizing a tile-based approach.
2.3 Training Data for High-resolution Depth Estimation
High-resolution DE requires datasets with fine-grained depth details and minimal label noise. Real-world datasets [12, 37, 4, 15, 22] are often sparse due to the limitations of LiDAR or contain inaccurate depth boundaries when obtained through Structure-from-Motion (SfM). Therefore, previous researches [7, 26] have trained the networks using a small number of relatively dense real-world datasets [44, 20, 34]. Some studies [7, 21] utilize pseudo labels for self-supervision, but these labels often contain noise. Other works [23, 24] leverage synthetic datasets for more accurate depth, resulting in degenerate generalization abilities due to the domain gaps. To take advantage of dense depth annotations in synthetic datasets without compromising generalization performance, we introduce Bias Free Masking (BFM), which selectively identifies the regions where supervision is applied. It mitigates overfitting to synthetic dataset biases and enhances robustness in zero-shot high-resolution DE.
3 Method
Figure 3: Labeling error examples in UnrealStereo4K [38] dataset.
Figure 4: Overview of the Framework. (a) Patch-wise Refinement Process. First, the pretrained MDE model $\Psi$ estimates a coarse depth $\mathbf{D}_{\mathrm{c}}$ and a fine depth $\mathbf{D}_{\mathrm{f}}^{i}$ from the resized input image $\mathbf{I}$ and the $i$-th patch $\mathbf{P}^{i}$. Based on $\mathbf{D}_{\mathrm{c}}$, the residual prediction network predicts residuals $\mathbf{R}$. This process is applied to every patch. Note that Bias Free Masking (BFM) is only required at training. (b) Grouped Patch Consistency Training. The training image is cropped with overlapping regions, followed by patch-wise refinement. Subsequently, depth consistency loss $\mathcal{L}_{\mathrm{con}}$ is applied to the gray-shaded regions to enforce consistency between depth results refined separately. (c) Bias Free Masking. We identify the reliable region as $\mathbf{M}_{\mathrm{BFM}}$ by utilizing $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$.
We describe our Patch Refine Once (PRO) model. In Sec. 3.1, we explain the process of PRO. Then, Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM), which are the key contributions of our paper, are described in Sec. 3.2 and Sec. 3.3, respectively.
3.1 Overview of Patch Refine Once (PRO)
Fig. 4 illustrates the overall pipeline of the proposed PRO framework. Following [23], given an original input image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, we estimate a coarse depth map $\mathbf{D}_{\mathrm{c}}$ by feeding a resized version of $\mathbf{I}$ into a pretrained zero-shot MDE network $\Psi$. $\mathbf{D}_{\mathrm{c}}$ captures the overall scene structure, which is essential for preserving global consistency.
To complement the local details, we take the following steps. First, we crop the original image $\mathbf{I}$ into $N$ grid-based patches $\{\mathbf{P}^{i}\}_{i=1}^{N}$. Each individual patch $\mathbf{P}^{i}$ is then resized and fed into $\Psi$ to estimate the corresponding fine depth map $\mathbf{D}_{\mathrm{f}}^{i}$. We extract two sets of five-level features from the decoder of $\Psi$: (i) $\mathbf{F}_{\mathrm{c}}=\{\mathbf{f}_{\mathrm{c,j}}^{i}\}_{j=1}^{5}$ for the resized image and (ii) $\mathbf{F}_{\mathrm{f}}=\{\mathbf{f}_{\mathrm{f},j}^{i}\}_{j=1}^{5}$ for $\mathbf{P}^{i}$. To align $\mathbf{F}_{\mathrm{c}}$ with $\mathbf{F}_{\mathrm{f}}$, we apply the $\mathsf{ROI}$ operation [16] to extract patch-aligned features from $\mathbf{F}_{\mathrm{c}}$. The patch-level features $\mathbf{F}_{\mathrm{f}}^{i}$ and the corresponding $\mathsf{ROI}$-extracted features are subsequently passed into a fusion module within the residual prediction network $\theta$. The fusion module integrates these features using a wavelet transform [8] and shallow convolutional layers, effectively combining global and local information. Finally, the refined depth map for $\mathbf{P}^{i}$, denoted as $\mathbf{D}_{\mathrm{refine}}^{i}$, is obtained as:
$\mathbf{D}_{\mathrm{refine}}^{i}=\theta(\mathsf{concat}(\mathbf{P}^{i},\mathsf{ROI}(\mathbf{D}_{\mathrm{c}}),\mathbf{D}_{\mathrm{f}}^{i})),$ (1)
where $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}})=\mathbf{D}_{\mathrm{c}}^{i}$ is a region of interest extraction operator for $\mathbf{D}_{\mathrm{c}}$ for $\mathbf{P}^{i}$, yielding $\mathbf{D}_{\mathrm{c}}^{i}$, and $\mathsf{concat}$ denotes channel-wise concatenation. Unlike previous methods [23, 24] where separate networks are trained to predict $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{f}}$ individually, we utilize the same pretrained zero-shot MDE network $\Psi$ and train only the residual prediction network as shown in Fig. 4-(a).
3.2 Grouped Patch Consistency Training (GPCT)
To address the boundary artifacts introduced by patch-wise refinement, we propose a simple yet effective GPCT strategy that ensures depth consistency across patch boundaries. While previous approaches [23, 24] focus only on two diagonally adjacent overlapping patches (e.g., (A, D) or (B, C) in Fig. 4) to enforce consistency constraints, our method employs four overlapping patches (A, B, C, and D in Fig. 4) simultaneously. As shown in Fig. 4-(b), we divide the training sample into overlapping patches, so that each patch overlaps with its neighbors. After refining each patch independently, we apply a depth consistency loss $\mathcal{L}_{\text{con}}$ to enforce consistency between the depth refinement results of overlapping patches:
$\mathcal{L}_{\text{con}}=\sum_{i\neq j}\frac{1}{|\Omega|}\sum_{p\in\Omega}\left(\mathbf{D}_{\mathrm{refine}}^{i}(p)-\mathbf{D}_{\mathrm{refine}}^{j}(p)\right)^{2},$ (2)
where $\mathbf{D}_{\mathrm{refine}}^{i}$ and $\mathbf{D}_{\mathrm{refine}}^{j}$ denote the depth predictions from overlapping patches $i$ and $j$, respectively. We denote the overlapping region between patches $i$ and $j$ as $\Omega$, with $|\Omega|$ indicating the number of pixels within this region. The merged depth map $\mathbf{D}_{\text{merged}}$ at position $p$ is computed as:
$\mathbf{D}_{\mathrm{merged}}(p)=\frac{1}{N_{o}(p)}\sum_{i\in\{A,B,C,D\}}\mathbf{D}_{\mathrm{refine}}^{i}(p),$ (3)
where $\mathbf{D}_{\mathrm{refine}}^{i}$ is the predicted depth from patch $i$ where $i=A,B,C$ and $D$ shown in Fig.4-(b), and $N_{o}(p)$ is the number of overlapping patches contributing depth estimates at location $p$. Then, we calculate the loss between merged depth $\mathbf{D}_{\mathrm{merged}}$ and GT depth $\mathbf{D}_{\mathrm{gt}}$. Since our method computes the loss using all four patches simultaneously, every backward propagation step enforces consistency at all boundaries. In contrast, the prior method [23] only applied a consistency loss on the overlapped region between two patches (e.g., (A, D) or (B, C)) at a time, leading to insufficient boundary supervision. Our approach provides a stronger supervision signal, yielding significantly improved cross-patch consistency and smooth boundary transitions during inference. By training with this GPCT strategy, our PRO model mitigates boundary artifacts without requiring additional refinement during inference, allowing efficient patch-wise depth estimation, as only a single refinement step per patch (i.e., “One Look”) is required.
3.3 Bias Free Masking (BFM)
To utilize dense GT depth when training depth refinement models for fine-grained details, we use a synthetic dataset such as the UnrealStereo4K [38]. However, the UnrealStereo4K dataset, which provides 4K resolutions, annotates the depths of transparent objects as the depths of their backgrounds (e.g., window objects in Fig. 3), which becomes a specific bias to the dataset. To maintain the benefits of dense depth labeling while avoiding overfitting to the dataset-specific biases, we propose Bias Free Masking (BFM) that uses the prior knowledge of the pretrained zero-shot MDE models. That is, the strategy of BFM is to exclude the unreliable regions corresponding to the incorrect GT depths for the transparent objects in the synthetic dataset (UnrealStereo4K) during training. Since the pretrained zero-shot MDE model $\Psi$ can be considered to have informative prior knowledge learned from large scale image-depth pairs, they can be used to identify the unreliable regions where $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ might significantly deviate each other, indicating potential biases to the synthetic dataset.
  Methods Publications Runtime (s) Booster ETH3D Middle14 NuScnes DIS
[…]
DepthAnythingV2 NIPS 2024 - 0.0307 0.993 0.0465 0.983 0.0307 0.994 0.0979 0.106 0.882 0.059
BoostingDepth CVPR 2021 12.7 0.0330 0.993 0.0552 0.974 0.0330 0.995 0.1035 0.115 0.870 0.170
PatchFusion P=16 CVPR 2024 3.4 0.0504 0.985 0.0735 0.956 0.0450 0.989 0.1124 0.141 0.831 0.206
PatchFusion P=177 CVPR 2024 36.8 0.0496 0.986 0.0723 0.957 0.0448 0.989 0.1050 0.139 0.833 0.189
PatchRefinerP=16 ECCV 2024 1.5 0.0348 0.989 0.0435 0.985 0.0292 0.995 0.0830 0.107 0.879 0.151
PatchRefinerP=177 ECCV 2024 16.2 0.0336 0.991 0.0430 0.985 0.0292 0.995 0.0805 0.106 0.881 0.141
PRO (Ours) - 1.4 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.104 0.883 0.156
Table 1: Quantitative comparison of depth estimation methods on Booster [29], ETH3D [35], Middlebury 2014 [34], NuScenes [4], and DIS-5K [28] datasets. Bold indicates the best performance in each metric.
Given a resized version for input image $\mathbf{I}$, we obtain coarse depth $\mathbf{D}_{\mathrm{c}}$ from $\Psi$ and GT depth $\mathbf{D}_{\mathrm{gt}}$ from the synthetic dataset. Then, by measuring the relative consistency between $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$, we identify unreliable regions as:
$\mathbf{M}_{\mathrm{unreliable}}=\left[\max\left(\frac{N(\mathbf{D}_{\mathrm{c}})}{N(\mathbf{D}_{\mathrm{gt}})},\frac{N(\mathbf{D}_{\mathrm{gt}})}{N(\mathbf{D}_{\mathrm{c}})}\right)>\tau\right],$ (4)
[…]
$\mathbf{E}_{\mathrm{c}}=\text{dilate}(\text{edge}(\mathbf{D}_{\mathrm{c}})),\quad\mathbf{E}_{\mathrm{gt}}=\text{dilate}(\text{edge}(\mathbf{D}_{\mathrm{gt}})),$ (5)
to guarantee that the final edge mask should not miss the regions that require depth refinements, otherwise they might be excluded simply from complementing $\mathbf{M}_{\mathrm{unreliable}}$. We define the edge mask as $\mathbf{M}_{\mathrm{edge}}=\mathbf{E}_{\mathrm{c}}\cap\mathbf{E}_{\mathrm{gt}}$. Finally, the BFM that denotes the reliable mask is constructed as $\mathbf{M}_{\mathrm{BFM}}=\textbf{M}_{\mathrm{edge}}\cup\sim\textbf{M}_{\mathrm{unreliable}}$ where $\sim\textbf{M}_{\mathrm{unreliable}}$ is the complement of $\textbf{M}_{\mathrm{unreliable}}$. This ensures that the unreliable regions due to large discrepancies between $\mathbf{D}_{\mathrm{c}}$ and $\mathbf{D}_{\mathrm{gt}}$ can be ignored, while preserving essential edge details for training. Consequently, the supervision can only be applied for the reliable regions via a masked loss:
$\mathcal{L}_{\mathrm{masked}}=\frac{1}{|\textbf{M}_{\mathrm{BFM}}|}\sum_{p\in\textbf{M}_{\mathrm{BFM}}}\mathcal{L}(\mathbf{D}_{\mathrm{merged}}(p),\mathbf{D}_{\mathrm{gt}}(p))$ (6)
where $\mathcal{L}$ is the loss between merged depth and GT depth, which is the combination of an L1 loss, an L2 loss, and a multi-scale gradient loss with four scale levels [22] using a ratio of 1:1:5. By restricting training to reliable regions, the depth refinement model refines depth only where necessary, avoiding modifications in regions where the pretrained model’s prior knowledge is considered more reliable than the (noisy) GT. This helps prevent the model from refining transparent regions incorrectly.
Final Loss function. The final training loss consists of masked loss $\mathcal{L}_{\mathrm{masked}}$ and depth consistency loss $\mathcal{L}_{\mathrm{con}}$ as:
$\mathcal{L}_{\mathrm{final}}=\mathcal{L}_{\mathrm{masked}}+\lambda\mathcal{L}_{\mathrm{con}}.$ (7)
where $\lambda$ is empirically set to 4 in our experiments.
4 Experiments
4.1 Datasets and Metrics
We train recent SOTA depth refinement models and our PRO model on a synthetic dataset to get the advantage of dense GT depth for a fair comparison.
UnrealStereo4K. The UnrealStereo4K dataset [38] is a synthetic dataset comprising of stereo images with 4K resolution (2,160$\times$3,840) and pixel-wise GT disparity labels. As aforementioned, its some GT labels are inaccurate especially for transparent objects as shown in Fig. 3.
Booster. The Booster dataset [29] has high-resolution (3,008 $\times$ 4,112) indoor images with specular and transparent surfaces. We use the training set with GT depth for testing. To ensure diversity, we removed images captured from the same scenes, resulting in a test set of 38 images.
ETH3D. The ETH3D dataset [35] contains high-resolution indoor and outdoor images (6,048 $\times$ 4,032) with GT depth captured by LiDAR sensors. Among the dataset, 34 images were rotated 90 degrees counterclockwise, causing the bottom sides to appear on the right sides. Before testing, we rotate these images 90 degrees clockwise to ensure that the floor sides are correctly positioned at the bottom sides.
Middlebury 2014. The Middlebury 2014 dataset [34] consists of 23 high-resolution (nearly 4K) indoor images with dense ground truth disparity maps.
NuScenes. The NuScenes dataset [4] is an autonomous driving dataset that provides multi-directional camera images along with depth information obtained from LiDAR and radar sensors. Among outdoor datasets, the Cityscapes dataset [6] provides higher resolution but suffers from blurred edges since its disparity maps are generated using a stereo matching algorithm instead of LiDAR sensors. Therefore, we use the NuScenes dataset for testing.
DIS-5K. DIS-5K dataset [28] is a high-resolution image segmentation dataset that provides accurately annotated masks, making it suitable for evaluating edge accuracy.
Evaluation Metrics. We use the commonly adopted metrics for DE, Absolute Relative Error (AbsRel) and $\delta_{1}$. Additionally, to evaluate the edge quality, we adopt $\mathrm{D}^{3}\mathrm{R}$ metric on the Middlebury 2014 dataset follwing [26], and measure Boundary Recall (BR) metric on the DIS-5K dataset following [3]. We utilize Consistency Error (CE) [23] to evaluate the consistency of the DE results across patches.
Figure 5: Qulitative comparisons for patch-wise DE methods on ETH3D [35] and DIS-5k [28]. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
  Metric PatchFusion PatchRefiner PRO (Ours)
[…]
Table 2: Comparison of PatchFusion [23], PatchRefiner [24], and PRO (Ours) in terms of Consistency Error (CE).
[…]
(a) 0.0385 0.987 0.0428 0.985 0.0292 0.995 0.0807 0.105 0.882 0.208
(b) 0.0303 0.994 0.0426 0.985 0.0290 0.995 0.0837 0.105 0.882 0.117
(c) 0.0313 0.993 0.0425 0.985 0.0288 0.996 0.0792 0.105 0.882 0.058
(d) 0.0304 0.994 0.0422 0.985 0.0287 0.996 0.0803 0.104 0.883 0.049
Table 3: Ablation on GPCT and BFM of our PRO for Booster [29], ETH3D [35], Middlebury 2014 [34], and NuScenes [4] datasets.
4.2 Implementation Details
We adopt the pretrained DepthAnythingV2 (DA2) [47] as the baseline model ($\Psi$ in Fig. 4). PRO, PatchFusion [23], and PatchRefiner [24] are (re)trained based on DA2, while BoostingDepth [26] uses DA2 without additional training. The input resolution to $\Psi$ is fixed at $518\times 518$. When we use the GPCT strategy, adjacent patches are cropped with an overlap of 224 pixels, ensuring an overlap of $224\times h$ for horizontally adjacent patches and $224\times w$ for vertically adjacent patches, where $h$ and $w$ denote the height and width of the patch, respectively. In our BFM, the dilation kernel size is empirically set to (10, 20). Training is conducted on 7,592 UnrealStereo4K samples for 8 epochs with a batch size of 64 (with gradient accumulation), taking 10 hours on a single RTX 4090 GPU.
4.3 Performance Comparison
At the test time, each input image is divided into a $4\times 4$ grid of patches. The depth of each patch is refined individually, and the refined patches are then reassembled to generate the final depth map.
Zero-shot performance on high-resolution datasets. We evaluate four patch-wise DE models: three are depth refinement models including our PRO model, BoostingDepth [26], PatchRefiner [24], and the other one is a direct DE model that is PatchFusion [23]. For PatchFusion and PatchRefiner, Our evaluations are done with two conditions: (i) ‘$\mathrm{P}$=16’ setting that uses the same patch number as our method, and (ii) ‘$\mathrm{P}$=177’ setting, where additional patches are used for test-time ensembling. Note that $\mathrm{P}$ denotes the number of patches used to reassemble final depth maps during inference. As shown in Table 1, our PRO achieves SOTA performance across all depth metrics while maintaining the lowest inference time. Notably, our BFM prevents overfitting to synthetic dataset biases, leading to a 9.5% improvement in AbsRel on the Booster dataset that contains transparent objects. In terms of edge quality, our method performs worse than BoostingDepth and PatchFusion in Boundary Recall (BR). However, as shown in the depth metrics, these two methods prioritize enhancing edge sharpness rather than focusing on enhancing overall depth accuracy. In contrast, when compared to PatchRefiner that exhibits similar depth performance, our PRO achieves superior results in terms of the edge metric (BR). In terms of consistency, our PRO achieves the best result, with an 85.9% improvement, as shown in Table 2. It demonstrates the effectiveness of our GPCT strategy in maintaining consistency between individually processed patches.
Qualitative Comparisons. Fig. 5 shows qualitative comparisons for four patch-wise DE models. As shown in Fig. 5, PatchFusion and PatchRefiner trained on the synthetic dataset without masking exhibit artifacts in the window regions. Additionally, two patch-based methods [24, 23] produce noticeable boundary artifacts when test ensembling is not applied. In contrast, despite refining each patch only once, our PRO demonstrates minimal depth discontinuity while achieving the fastest inference time.
4.4 Ablation Studies
We analyze the effectiveness of the core components of our PRO through ablation studies: Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM).
Ablation on GPCT and BFM. Table 3 presents the ablation results on GPCT and BFM. When BFM is applied to the baseline (a), the improvements are relatively minor for most datasets. However, for the Booster dataset [29] that contains transparent objects, we observe a significant improvement of 21.3% in AbsRel metric, demonstrating that our BFM method effectively identifies unreliable regions. Even when only BFM is applied without GPCT, a reduction in consistency error (CE) suggests that supervision limited to reliable regions helps mitigate unnecessary refinement, thereby reducing confusion in the depth refinement process. Furthermore, applying GPCT to the baseline yields a 50.4% improvement in consistency error, demonstrating its effectiveness in maintaining consistency between independently processed patches. Although model (c) achieves the best performance in the $\mathrm{D}^{3}\mathrm{R}$ metric, it introduces artifacts in transparent regions. Consequently, model (d), which incorporates both GPCT and BFM, achieves the best overall performance across most metrics.
[…]
28 0.108 0.0423 0.0288
56 0.113 0.0424 0.0289
112 0.065 0.0425 0.0288
224 0.049 0.0422 0.0287
448 0.060 0.0423 0.0287
Table 4: Ablations on overlap sizes in Consistency Error (CE) for ETH3D [35] and Middlebury 2014 [34] datasets.
Ablation on Overlap Sizes in GPCT. Table 4 shows the effect of different overlap sizes in GPCT. Each overlap value represents the shorter side of the overlapping region between adjacent patches. As shown, increasing the overlap generally reduces the CE. However, when the overlap reaches to 448 pixels, CE begins to increase. We attribute this to the fact that highly overlapped patches capture nearly identical scene contents, resulting in fewer inconsistencies between patches. So, the depth consistency loss provides less meaningful supervision, reducing its effectiveness.
5 Conclusion
In this paper, we propose the Patch Refine Once (PRO) model for depth refinement on high-resolution images. To address the depth discontinuity problem, we introduce the Grouped Patch Consistency Training (GPCT) strategy, ensuring consistency between independently processed patches. Also, we propose Bias Free Masking (BFM) to prevent the depth refinement model from overfitting to dataset-specific biases. Through these strategies, our PRO achieves superior zero-shot performance while maintaining a low inference time (12$\times$ faster than PatchRefiner with 177 patches), outperforming recent state-of-the-art methods across diverse datasets.
APPENDIX
{strip}[Uncaptioned image]
Figure 6: Architecture of the Residual Prediction Network and Frequency Fusion Module (FFM). (a) Residual Prediction Network. The Residual Prediction Network comprises an encoder, a decoder, and the Fusion Module. (b) Frequency Fusion Module (FFM). We utilize Discrete Wavelet Transform (DWT) to decompose the input features into four frequency components. Then, each frequency component is processed independently using convolutional layers.
  Models FLOPs (Mac) Parameter DIS UHRSD Middle14 ETH3D Booster NuScenes
[…]
Conv 31664G 63M 0.127 0.071 0.0290 0.995 0.0842 0.0428 0.984 0.0306 0.993 0.106 0.882
FFM (Ours) 20250G 51M 0.156 0.083 0.0287 0.996 0.0803 0.0422 0.985 0.0304 0.994 0.104 0.883
Table 5: Ablation study of the Frequency Fusion Module (FFM) on DIS-5K [28], UHRSD [45], Middlebury 2014 [34], ETH3D [35], Booster [29], and NuScenes [4]. Bold indicates the best performance in each metric.
Appendix A Architecture of the Fusion Module
The encoder takes $\mathbf{P}^{i}$, $\mathbf{D}_{\mathrm{f}}^{i}$, and $\mathsf{ROI}(\mathbf{D}_{\mathrm{c}})$ as inputs and produces a set of five intermediate features, denoted as $\mathbf{F}_{\mathrm{enc}}=\{\mathbf{f}_{\mathrm{enc},j}^{i}\}_{j=1}^{5}$ following [23]. Then, the fused feature map $\mathbf{F}_{\mathrm{fuse}}$ is obtained through the frequency fusion module ($\mathsf{FFM}$), defined as $\mathbf{F}_{\mathrm{fuse}}=\mathsf{FFM}(\mathsf{concat}(\mathbf{F}_{\mathrm{f}},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}}))).$
Subsequently, $\mathsf{concat}(\mathbf{F}_{\mathrm{fuse}},\mathsf{ROI}(\mathbf{F}_{\mathrm{c}}),\mathbf{F}_{\mathrm{enc}})$ is processed through two consecutive layers, each consisting of a $3\times 3$ convolution, batch normalization, and ReLU activation. Finally, the resulting feature is fed into the DPT decoder [31] to obtain the residual map $\mathbf{R}$.
Architecture of the Frequency Fusion Module (FFM) To obtain accurate depth values from the coarse depth and preserve fine details from the fine depth, we design a Frequency Fusion Module (FFM) that effectively extracts and integrates edge information. We utilize Discrete Wavelet Transform (DWT) to decompose the input features into four frequency components: LL, LH, HL, and HH, which represent the low-frequency and high-frequency information. Each component is processed with its own dedicated convolution to capture scale-specific features. Finally, the components are recombined using the Inverse Discrete Wavelet Transform (IDWT), resulting in fused features that retain both global depth consistency and enhanced edge details. Overall process is described in Fig. 6-(b). To describe this process in more detail, we first decompose $\mathsf{ROI}(\mathbf{F}_{\mathrm{c}})$ and $\mathbf{F}_{\mathrm{f}}$ into four frequency sub-bands $\mathbf{X}$ ($\forall\mathbf{X}\in\{\mathbf{LL},\mathbf{LH},\mathbf{HL},\mathbf{HH}\}$) using $\mathsf{DWT}$. Each sub-band is then fused using a corresponding convolution $\text{Conv}_{\mathbf{X}}$. Finally, the fused feature map $\mathbf{F}_{\mathrm{fuse}}$ is obtained through $\mathsf{IDWT}$.
$\mathbf{X}_{\mathrm{c}},\mathbf{X}_{\mathrm{f}}=\mathsf{DWT}(\mathsf{ROI}(\mathbf{F}_{\mathrm{c}})),\mathsf{DWT}(\mathbf{F}_{\mathrm{f}})$ (8)
$\mathbf{X}_{\text{fuse}}=\text{Conv}_{\mathbf{X}}(\mathsf{concat}(\mathbf{X}_{\mathrm{c}},\mathbf{X}_{\mathrm{f}}))$ (9)
$\mathbf{F}_{\text{fuse}}=\mathsf{IDWT}(\mathbf{LL}_{\text{fuse}},\mathbf{LH}_{\text{fuse}},\mathbf{HL}_{\text{fuse}},\mathbf{HH}_{\text{fuse}})$ (10)
A.1 Ablation study of FFM
To validate the effectiveness of the Frequency Fusion Module (FFM), we conduct an ablation study by replacing the FFM with a simple convolutional block consisting of Conv-ReLU-Conv layers. To ensure that any performance gain is not simply due to an increase in the number of parameters or FLOPs, we design the simple convolutional block to have more parameters and FLOPs than the FFM. This allows us to attribute the performance improvement to the design of FFM itself, rather than computational complexity. As shown in the Table 5, our method not only achieves the best performance in standard depth metrics, but also yields significant improvements in edge accuracy. Specifically, it achieves a 22.5% improvement on the DIS-5K dataset and a 16.9% improvement on the UHRSD dataset in the Boundary Recall (BR) metric. In addition, we observe a 4.6% improvement in the edge quality metric (D3R). It demonstrates that the proposed FFM effectively integrates edge information through the use of Discrete Wavelet Transform (DWT), which enables selective enhancement of high-frequency details without sacrificing global structure. This highlights the benefit of frequency-domain processing in depth refinement tasks.
Appendix B Qualitative Results
Ablation Study In the ablation study (Sec. 4.4), we analyze the effect of Grouped Patch Consistency Training (GPCT) and Bias Free Masking (BFM) quantitatively. In this section, we analyze the effect of GPCT and BFM with qualitative results. As shown in the Fig. 7, the model trained without GPCT shows remarkable depth discontinuity problem on the grids. On the other hand, our PRO model trained with GPCT alleviates the depth discontinuity problem. Likewise, as shown in the Fig. 8 , the model trained without BFM exhibits artifacts on transparent surfaces, such as glass windows, as well as reflective surfaces like TV screens. In contrast, our model trained with BFM effectively refines only the edge regions while preventing artifacts on transparent objects.
Additional Qualitative Results
We provide additional qualitative comparisons of BoostingDepth [26], PatchFusion [23], PatchRefiner [24], and PRO (Ours) on the UHRSD [45] dataset and on internet images (e.g., from Unsplash [^1] https://unsplash.com and Pexels [^2] https://www.pexels.com), as shown in Fig.9 and Fig.10.
Figure 7: Qulitative comparisons of GPCT’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Grouped Patch Consistency Training (GPCT) (b). (b) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
Figure 8: Qulitative comparisons of BFM’s impact on ETH3D [35], Booster [29], and DIS-5k [28]. We compare PRO (Ours (d)) with the model trained without Bias Free Masking (BFM) (c). (c) and (d) represent the model index in Table 3. We also visualize the residuals to highlight the presence of artifacts more effectively. Zoom in for details.
Figure 9: Qulitative comparisons for patch-wise DE methods on UHRSD [45] and images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the leftmost depth column represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.
Figure 10: Qulitative comparisons for patch-wise DE methods on images from the internet. We compare PRO (Ours) with BoostingDepth [26], PatchFusion (PF) [23], and PatchRefiner (PR) [24]. The time displayed in the top depth row represents the inference time. Black rectangle area represents the transparent object region. Black arrows indicate patch boundaries. Zoom in for details.